Hura Corp. Has 500 convertible bonds outstanding. the bonds have a coupon rate of 5.5%, seven years to maturity and are currently priced at &1,250. The bonds have a conversion ratio of 25 and Hura Corp.'s common stock is currently priced at $52.35 per share. How much would you make per share (in $) by converting these bonds?
Answer #1

conversion ratio = 25

current price of the bond = 1250

so if we convert the bond now we get 25 shares

value of per share should be = 1250 / 25 = $50

current stock price = $52.35

so , gain per share = 52.35 - 50 = $2.35